# Misaligned incentives in the post-2024 UEFA Champions League: A managerial failure

### v6

@@ Abstract
UEFA declares that it is fully committed to respecting the fundamental values of sport. However, the qualification for the post-2024 UEFA Champions League is not always based on sporting merit as we show. In particular, a match with misaligned incentives has barely been avoided in the 2023/24 German Bundesliga. Our mathematical model reveals that fairness can be guaranteed by filling all vacant slots via the domestic leagues and retaining the association constraint in the league phase draw. UEFA repeatedly commits the same mistake in treating vacancies, which indicates a clear managerial failure. Therefore, governing bodies in sports are called to work more closely together with the scientific community.
Keywords: fairness; incentive compatibility; integrity; tournament design; UEFA Champions League
@@ 1 Introduction
Ensuring the fairness and integrity of sports is a crucial responsibility of tournament organisers (Kendall and Lenten,, 2017; Devriesere et al.,, 2024). Naturally, the Union of European Football Associations (UEFA) is also committed to these values, at least in declarations and media releases. For example, UEFA president Aleksander Čeferin has commented on the novel format of the UEFA Champions League in 2024 with the following words (UEFA, 2024a, ): “UEFA has clearly shown that we are fully committed to respecting the fundamental values of sport and to defending the key principle of open competitions, with qualification based on sporting merit, fully in line with the values and solidarity-based European sports model.
@@
The current paper uncovers that, while UEFA may be committed to sporting merit in principles, their regulations do not exclude an unfair game. In particular, although the recent fundamental reform of the UEFA Champions League has restored the integrity of the seeding system, the way in which the novel European performance spots is assigned allows for a situation when a team is strictly better off by losing. These misaligned incentives almost occurred in the last round of the 2023/24 German Bundesliga in the match Eintracht Frankfurt vs. RB Leipzig. A mathematical model is presented to understand the issue and compare different allocation rules of vacant slots. Two possible solutions are offered to decrease the probability of such a problem or even entirely avoid it. It is also shown that the association constraint in the league phase draw has a role in improving fairness.
The novel case of incentive incompatibility is especially worrying because “the objective for a sport organisation should be to eliminate any game that raises potential unfairness ex-ante(Scelles et al.,, 2024, p. 535). Furthermore, the root of the problem resides in the allocation of vacant slots, which led to some instances of incentive incompatibility in the past. Making the same mistake repeatedly seems to indicate a clear managerial failure.
@@ 2 Related literature
Since tanking—the act of deliberately losing a game—is against the spirit of sports, incentive compatibility is an extensively discussed topic in the field of tournament design (Kendall and Lenten,, 2017; Csató,, 2021). Probably the most serious type of incentive incompatibility is when a team can be strictly better off by losing. According to our knowledge, such a possibility has been identified first by Dagaev and Sonin, (2018) in the UEFA Europa League. Several qualification tournaments for the FIFA World Cups and the UEFA European Championships have been verified to suffer from the same weakness (Csató, 2020a, ), even allowing for a match where the dominant strategy of both opponents is to play a draw (Csató, 2020c, ).
@@
The introduction of the UEFA Nations League has also allowed for matches where a team has an incentive to lose (Csató,, 2022; Haugen and Krumer,, 2021). Even though Scelles et al., (2024) identify no such games between 2018 and 2021, the authors call for more research around optimal competition formats that provide appropriate incentives for all teams in all games. Our paper can be considered as one of these follow-up studies.
@@ 3 A real-world example
The four highest-ranked associations—including Germany—are currently represented by four clubs in the UEFA Champions League. In addition, both the UEFA Champions League and UEFA Europa League titleholders automatically qualify for the league phase. The possible vacancy is filled according to UEFA, 2024b (, Article 3.04):
@@
UEFA has introduced the so-called European performance spot from the 2024/25 season (UEFA, 2024b, , Article 3.07):
@@ 7 SC Freiburg 11 9 13 44 56 --12 42
Table 1 shows the standing of the 2023/24 German Bundesliga on the morning of 18 May 2024 (the last matchday) with a slight modification: the result of the match Borussia Dortmund vs. 1899 Hoffenheim, played on 15 February 2024, is changed from 2-3 to 4-3. At that moment, it was already known that Germany and Italy finished in the top two of the 2023/24 season association club coefficients, and earned the two European performance spots. The two clubs playing in the final of the UEFA Champions League on 1 June 2024 were known to be Borussia Dortmund and the Spanish club Real Madrid, too.
@@
Consider the situation of Eintracht Frankfurt, which plays against RB Leipzig at home on the last (34th) matchday. It could be ranked neither higher nor lower than its current sixth position. Since the top four teams automatically qualify for the UEFA Champions League, the uncertainty with respect to this tournament resides in whether Borussia Dortmund or RB Leipzig obtains the fourth position, and whether Borussia Dortmund wins the 2023/24 UEFA Champions League or not. They determine the fate of Eintracht Frankfurt as follows:
@@
  • If Borussia Dortmund is the fourth, then the first five teams qualify for the UEFA Champions League independently of which team wins the 2023/24 UEFA Champions League final. If Dortmund wins the final, a vacancy is created, filled by the Ukrainian champion Shakhtar Donetsk (UEFA, 2024b, , Article 3.04). If Dortmund loses the final, this team creates no vacancy.
@@
  • If Borussia Dortmund is the fifth, then the first four teams automatically qualify for the UEFA Champions League. If Dortmund wins the 2023/24 UEFA Champions League final, it qualifies for the next season of this tournament as the titleholder. Thus, the European performance spot is obtained by the sixth-ranked Eintracht Frankfurt. If Dortmund loses the final, it qualifies for the next season of the UEFA Champions League due to the European performance spot earned by the Bundesliga.
@@
In the first case, Eintracht Frankfurt could never play in the UEFA Champions League. However, in the second case, it can participate in this competition with a non-negligible probability (if Dortmund wins the 2023/24 UEFA Champions League). Consequently, the dominant strategy of Eintracht Frankfurt is to guarantee that Borussia Dortmund is ranked fifth in the 2023/24 German Bundesliga, which can be easily achieved by deliberately losing against RB Leipzig on 18 May 2024.
@@ 4 A mathematical model
The allocation rule is simplified if the order of Articles 3.04 and 3.07 in the UEFA Champions League regulation (UEFA, 2024b, ) is reversed. Then the two European performance spots are allocated first (UEFA, 2024b, , Article 3.07), followed by addressing the possible vacancies created by the titleholders (UEFA, 2024b, , Article 3.04), which leads to:
@@
Now the incentive (in)compatibility of these allocation rules can be explored.
@@
  • $Q_{i}\left(S\right)=Q_{i}\left(S^{\prime}\right)$;
  • $Q_{j}\left(S\right)\geq 2$ but $Q_{j}\left(S^{\prime}\right)=1$ for a team $1\leq j\leq\lvert S\rvert$.
@@
That is, a tanking strategy by team $i$ guarantees that it remains ranked $(q+1)$th, while reverses teams $j$ and $k$ in the two positions directly above.
In tournament $(U,R)$, no tanking strategy exists because the UEFA club coefficients cannot be increased by a worse performance on the field.
@@
If the third condition is satisfied, team $i$ can create a vacancy by achieving that the winner of tournament $(S,P)$ is ranked $q$th in tournament $(T,Q)$, which is filled by team $i$.
The definition of the allocation rule $\mathcal{A}$ implies that no other team can receive the possible vacant slot, thus, incentive incompatibility can arise only in the cases above. ∎
The example presented in Section 3 exploits the third condition.
@@ 5 Discussion
A recent survey on fraud in sport (Vanwersch et al.,, 2022) concludes that future research is needed to accumulate more knowledge on the prevention of fraud. Our paper examines the new format of the UEFA Champions League from this perspective. The rules have clearly improved from the 2024/25 season. Now the clubs are seeded based on their individual club coefficients except for the guaranteed place of the titleholder in the first pot (UEFA, 2024b, , Article 16.01). Therefore, the regime used until the 2014/15 season has returned, and the seeding system satisfies incentive compatibility.
@@
Consequently, even though rule change A seems to be closer to the original intention of UEFA, it might still lead to incentive incompatibility, albeit with a lower probability. At least, the situation presented in Section 3 will certainly be avoided. In contrast, rule change B entirely avoids misaligned incentives and highlights the role of the association constraint in the league phase draw, which is desirable not only to maximise the number of international matches played but also to ensure strategy-proofness in the league phase. This is a good illustration of the idea raised by Csató, (2022) that draw restrictions can be used to prohibit potentially unfair matches. Nonetheless, rule change B favours the domestic league of the titleholder, hence, it leads to a somewhat higher concentration of the clubs playing in the UEFA Champions League by their associations because the highest-ranked countries are more likely to win both the UEFA Champions League and the UEFA Europa League.
@@
If UEFA is not worried much about manipulation in a domestic league and is not willing to change its regulations, the national associations are also able to mitigate the problem to some extent. The current paper has considered incentive incompatibility as a binary concept. However, the probability of a situation that is vulnerable to tanking can be quantified via Monte Carlo situations, analogously to Csató, (2022), even if it would be difficult to implement.
Naturally, the set of matches to be played cannot be changed in a domestic championship without changing the tournament format, but the order of the matches is a decision variable. The scenario outlined in Section 3 has become threatening especially since the match Eintracht Frankfurt vs. RB Leipzig was played in the last round. The set of top teams may be relatively accurately estimated at the beginning of the season, thus, the last rounds can be scheduled to avoid any matches between the strongest teams. Such a schedule might lead to a tanking opportunity at the end of the season with a (much) lower probability. But even if Eintracht Frankfurt had already played against RB Leipzig, it would be unfair if it turns out ex post that a team has missed qualifying for the UEFA Champions League due to winning a match months before. Therefore, it is the responsibility of UEFA to act if it is really committed to its announced principles.
6 Conclusions
Our paper has revealed that the UEFA Champions League rules are incentive incompatible again from the 2024/25 season. In particular, the allocation rules of the vacancies caused by the titleholders and the two European performance spots have almost led to a game with misaligned incentives in the 2023/24 German Bundesliga. UEFA is strongly encouraged to adopt our recommendation in order to safeguard the integrity of its flagship tournament.

### v7

# Misaligned incentives in sports tournaments: A mathematical analysis of the post-2024 UEFA Champions League

@@ Abstract
UEFA declares that it is fully committed to respecting the fundamental values of sport. However, the rules of the post-2024 UEFA Champions League are shown to be unfair: a match in the 2023/24 German Bundesliga almost created powerful incentives for a team to lose. We develop a mathematical model to reveal how fairness can be ensured by giving priority to domestic leagues in filling vacancies. The implementation of our recommendations reduces or even eliminates any similar problem with misaligned incentives. It is also presented that UEFA repeatedly commits the same mistake by allocating the vacant slots unfairly. Therefore, governing bodies in sports are called to work more closely together with the scientific community.
Keywords: fairness; incentives; integrity; tournament design; UEFA Champions League
@@ 1 Introduction
Ensuring the fairness and integrity of sports is a crucial responsibility of tournament organisers (Forrest and McHale,, 2019; Kendall and Lenten,, 2017; Devriesere et al.,, 2024). One of the most obvious criteria of fairness is incentive compatibility, in other words, the lack of misaligned incentives: it should never happen that one agent (player/team) seeks to win and the other to lose. This would be clearly against the spirit of the game and may easily lead to collusion or match-fixing (Csató,, 2024; Haigh,, 2009).
Naturally, the Union of European Football Associations (UEFA) is also committed to these values, at least in declarations and media releases. For example, UEFA president Aleksander Čeferin has commented on the novel format of the UEFA Champions League in 2024 with the following words (UEFA, 2024a, ): “UEFA has clearly shown that we are fully committed to respecting the fundamental values of sport and to defending the key principle of open competitions, with qualification based on sporting merit, fully in line with the values and solidarity-based European sports model.
@@
The current paper uncovers that, while UEFA may be committed to sporting merit in principles, their regulations do not always exclude an unfair game with misaligned incentives. Even though the recent reform of the UEFA Champions League improved the seeding system, the assignment of the novel European performance spots may create an incentive to lose. This has almost occurred in the last round of the 2023/24 German Bundesliga in the match Eintracht Frankfurt vs. RB Leipzig. To analyse the issue in depth, we develop a mathematical model to propose two solutions that can improve fairness.
The novel case of incentive incompatibility is especially worrying because “the objective for a sport organisation should be to eliminate any game that raises potential unfairness ex-ante(Scelles et al.,, 2024, p. 535). Furthermore, the root of the problem resides in the allocation of vacant slots, which led to some instances of incentive incompatibility in the past. It seems that filling vacancies poses a significant challenge for tournament organisers, even though several academic papers have been published on this issue (see Section 2 and Table 2). Therefore, governing bodies in major sports, especially UEFA, are called to strengthen collaboration with the scientific community.
@@ 2 Related literature
Tanking—the act of deliberately losing a game—is against the spirit of sports and is a well-known violation of fairness in sports (Kendall and Lenten,, 2017; Csató,, 2021). Tanking is closely related to the axiom of incentive compatibility, which requires that tournament rules should encourage honest performance under any circumstances; otherwise, a team could face misaligned or perverse incentives that might inspire tanking.
Probably the most serious type of incentive incompatibility is when a team can be strictly better off by losing. According to our knowledge, such a possibility has been identified first by Dagaev and Sonin, (2018) in the UEFA Europa League. Several qualification tournaments for the FIFA World Cups and the UEFA European Championships have been verified to suffer from the same weakness (Csató, 2020a, ), even allowing for a match where the dominant strategy of both opponents is to play a draw (Csató, 2020c, ).
@@
The introduction of the UEFA Nations League has also allowed for matches where a team has an incentive to lose (Csató, 2022a, ; Haugen and Krumer,, 2021). Even though Scelles et al., (2024) identify no such games between 2018 and 2021, the authors call for more research around optimal competition formats that provide appropriate incentives for all teams in all games. Analogously, a recent overview of the causes and consequences of fraud in sport (Vanwersch et al.,, 2022) concludes that future research is necessary to help the competent authorities in tackling fraud and to gather more knowledge on actionable elements with regard to its prevention. Our paper can be considered as one of these follow-up studies.
@@ 3 A real-world example
The UEFA Champions League is the most prestigious European club football tournament. The number of participating teams for each national association is determined by the access list, which ranks the associations based on their UEFA association coefficients (Csató, 2022b, ). The clubs can enter the main league phase of the UEFA Champions League in several ways:
  • The highest-ranked associations have a given number of slots in the league phase. For instance, the four highest-ranked associations—including Germany—are currently represented by four clubs.
  • Teams from lower-ranked associations play in the UEFA Champions League qualification. The qualifying consists of two independent knockout competitions; the champions path for the champions of the lower-ranked associations and the league path for the teams of some higher-ranked associations in order to guarantee that a given number of champions play in the league phase.
  • Both the UEFA Champions League and UEFA Europa League (the second most prestigious European club football competition) titleholders automatically qualify for the league phase, even if they do not qualify for the league phase through their domestic championship.
  • The two associations with the most coefficient points in the previous season have one European performance spot each in the league phase.
However, if a titleholder qualifies for the league phase via its domestic championship, then a vacancy is created in the league phase, which is filled according to UEFA, 2024b (, Article 3.04):
@@
The so-called European performance spot has been introduced from the 2024/25 season (UEFA, 2024b, , Article 3.07):
@@ 7 SC Freiburg 11 9 13 44 56 --12 42
  • The first four teams automatically qualify for the UEFA Champions League due to the access list.
  • The fifth team also qualifies due to the European performance spot. However, if Dortmund is ranked fifth and wins the Champions League final, this slot will be obtained by the sixth-ranked team.
@@
Table 1 shows the standing of the 2023/24 German Bundesliga on the morning of 18 May 2024 (the last matchday) with a slight modification: the result of the match Borussia Dortmund vs. 1899 Hoffenheim, played on 15 February 2024, is changed from 2-3 to 4-3. At that moment, it was already known that Germany and Italy finished in the top two based on the 2023/24 season association club coefficients, and earned the two European performance spots. The two clubs playing in the final of the UEFA Champions League on 1 June 2024 were known to be Borussia Dortmund and the Spanish club Real Madrid, too.
@@
Consider the situation of Eintracht Frankfurt, which plays against RB Leipzig at home on the last (34th) matchday. It could be ranked neither higher nor lower than its current sixth position. Since the top four teams automatically qualify for the UEFA Champions League, the uncertainty with respect to this tournament resides in whether Borussia Dortmund or RB Leipzig obtains the fourth position, and whether Borussia Dortmund wins the 2023/24 UEFA Champions League or not.
The fate of Eintracht Frankfurt is determined as follows:
  • If Borussia Dortmund is the fourth and
    • $\diamond$
@@
      Dortmund wins the 2023/24 UEFA Champions League final, then a vacancy is created in the league phase, which is filled by the Ukrainian champion Shakhtar Donetsk according to UEFA, 2024b (, Article 3.04);
    • $\diamond$
      Dortmund loses the final, then a vacancy may be created by the UEFA Champions League titleholder Real Madrid, which would be filled by the Ukrainian champion Shakhtar Donetsk according to UEFA, 2024b (, Article 3.04).
    Consequently, the European performance spot obtained by the German Bundesliga is given to the fifth-ranked team RB Leipzig according to UEFA, 2024b (, Article 3.07). Thus, the first five teams qualify for the UEFA Champions League independently of which team wins the 2023/24 UEFA Champions League final.
  • If Borussia Dortmund is the fifth and
    • $\diamond$
@@
      Dortmund wins the 2023/24 UEFA Champions League final, then it qualifies for the next season of this tournament as the titleholder, and the European performance spot is given to the sixth-ranked Eintracht Frankfurt according to UEFA, 2024b (, Article 3.07);
    • $\diamond$
      Dortmund loses the final, then Dortmund obtains the European performance spot of the German Bundesliga according to UEFA, 2024b (, Article 3.07), while a vacancy may be created by the UEFA Champions League titleholder Real Madrid, which would be filled by the Ukrainian champion Shakhtar Donetsk according to UEFA, 2024b (, Article 3.04).
@@
In the first case when Dortmund is ranked fourth, Eintracht Frankfurt could never play in the UEFA Champions League. However, in the second case when Dortmund is ranked fifth, Eintracht Frankfurt can enter the Champions League with a non-negligible probability (if Dortmund wins the 2023/24 UEFA Champions League). Hence, the dominant strategy of Eintracht Frankfurt is to guarantee that Borussia Dortmund is ranked fifth in the 2023/24 German Bundesliga, which can be easily achieved by losing against RB Leipzig on 18 May 2024.
Is the team ranked between 1st and 4th?Does Borussia Dortmundwin the 2023/24 UEFA Champions League final?Is the 5th-ranked team Borussia Dortmund?Is the team ranked 5th?Is the team ranked 5th?Is the team ranked 6th?Qualification (access list)Qualification (EPS)Qualification (Titleholder)EliminationYesNoYesNoYesYesNoNoYesYesNo
Figure 1: Qualification for the 2024/25 UEFA Champions League
from the 2023/24 German Bundesliga
Figure 1 provides a graphical illustration of how the 2024/25 Champions League spots are allocated for the German teams. Obviously, the only way for the sixth-ranked team to qualify is if Borussia Dortmund is ranked fifth and wins the 2023/24 UEFA Champions League final simultaneously.
Note that the main problem resides in the strange interaction between UEFA, 2024b (, Article 3.04) and UEFA, 2024b (, Article 3.07). In particular, the order in which they are applied depends on the ranking in the 2023/24 German Bundesliga that can be exploited by a tanking strategy outlined in Dagaev and Sonin, (2018): in a round-robin tournament, a situation always exists in which a deliberate loss does not affect the rank of a particular team (in our case, Eintracht Frankfurt) but changes the order of two higher-ranked teams (in our case, RB Leipzig and Borussia Dortmund).
@@ 4 A mathematical model
This section presents a formal analysis of the problem. However, there is no need to understand the mathematical framework for practitioners in sports organizations or policymakers who are mainly interested in the implications since all these issues will be thoroughly discussed in Section 5.
@@
The allocation rule can be simplified if the order of Articles 3.04 and 3.07 in the UEFA Champions League regulation (UEFA, 2024b, ) is reversed. Then the two European performance spots are allocated first (UEFA, 2024b, , Article 3.07), followed by addressing the possible vacancies created by the titleholders (UEFA, 2024b, , Article 3.04), which leads to:
@@
Now the incentive (in)compatibility of these allocation rules will be explored.
@@
  • $P_{i}\left(S\right)=P_{i}\left(S^{\prime}\right)$;
  • $P_{j}\left(S\right)\geq 2$ but $P_{j}\left(S^{\prime}\right)=1$ for a team $1\leq j\leq\lvert S\rvert$.
@@
That is, a tanking strategy by team $i$ guarantees that it remains ranked $(q+1)$th, while reverses teams $j$ and $k$ in the two positions directly above ($(q-1)$th and $q$th).
In tournament $(U,R)$, no tanking strategy exists because the UEFA club coefficient cannot be increased by a worse performance on the field.
@@
If the third condition is satisfied, team $i$ can create a vacancy by ensuring that the winner of tournament $(S,P)$ is ranked $q$th in tournament $(T,Q)$, which is filled by team $i$.
The definition of the allocation rule $\mathcal{A}$ implies that no other team can receive the possible vacant slot, thus, incentive incompatibility arises only in the cases above. ∎
The first condition becomes relevant if a club (team $r$) that hopes to have the highest individual club coefficient in the champions path in the next season supports a club (team $j$) from a higher-ranked national association to win the UEFA Champions League in order to create a vacancy to be filled by itself. This is much less threatening than the third condition because tanking is excluded in the final knockout stage and the titleholder usually qualifies via its domestic championship, too. For example, the Ukrainian Shakhtar Donetsk (that had the highest coefficient in the champions path as we have seen in Section 3) played in a group with Antwerp, Barcelona, and Porto in the the 2023/24 Champions League. Therefore, it would be interested in supporting Barcelona to win the Champions League at the expense of Antwerp since it was more likely that Barcelona would also qualify through its domestic championship.
The second condition says that a club (team $i$) that is ranked $(q+1)$th in its domestic league supports the club (team $j$) ranked directly above it from the same league to win the UEFA Champions League in order to create a vacancy to be filled by itself. However, this cannot be beneficial because (1) tanking is excluded in the final knockout stage and (2) teams from the same national association are not allowed to play against each other in the previous (group or league) stage. The latter observation highlights the role of draw constraints with respect to incentive compatibility, which is extensively studied by Csató, 2022a . Thus, neither hypothetical nor historical cases could be constructed to bridge the gap between theory and practice.
The example presented in Section 3 exploits the third condition: $q=5$ implies the sixth-ranked Eintracht Frankfurt (team $i$) can benefit from swapping the fourth-ranked Borussia Dortmund (team $j$) and the fifth-ranked RB Leipzig (team $k$) if Dortmund plays in the Champions League and wins this competition.
@@ 5 Discussion
Recent papers on fairness, fraud, and incentives in sport (Scelles et al.,, 2024; Vanwersch et al.,, 2022) call for future research to accumulate more knowledge on this topic. Our paper examines the new format of the UEFA Champions League from the perspective of incentive compatibility, that is, whether the rules always provide the right incentives for the teams to perform. The rules have clearly improved from the 2024/25 season. Now the clubs are seeded based on their individual club coefficients except for the guaranteed place of the titleholder in the first pot (UEFA, 2024b, , Article 16.01). Therefore, the regime used until the 2014/15 season has returned, and the seeding system satisfies incentive compatibility.
@@
Consequently, even though rule change A seems to be closer to the original intention of UEFA, it might still lead to incentive incompatibility, albeit with a lower probability. At least, the situation presented in Section 3 will certainly be avoided. In contrast, rule change B entirely avoids misaligned incentives and highlights the role of the association constraint in the league phase draw, which is desirable not only to maximise the number of international matches played but also to ensure strategy-proofness in the league phase. This is a good illustration of the idea raised by Csató, 2022a that draw restrictions can be used to prohibit potentially unfair matches. Nonetheless, rule change B favours the domestic league of the titleholder, hence, it leads to a somewhat higher concentration of the clubs playing in the UEFA Champions League by their associations because the highest-ranked countries are more likely to win both the UEFA Champions League and the UEFA Europa League.
@@
If UEFA is not worried much about manipulation in a domestic league and is not willing to change its regulations, the national associations are also able to mitigate the problem to some extent. The current paper has considered incentive incompatibility as a binary concept. However, the probability of a situation that is vulnerable to tanking can be quantified via Monte Carlo situations, analogously to Csató, 2022a , even if it would be difficult to implement.
Naturally, the set of matches to be played cannot be changed in a domestic championship without changing the tournament format, but the order of the matches is a decision variable. The scenario outlined in Section 3 has become threatening especially since the match Eintracht Frankfurt vs. RB Leipzig was played in the last round. The set of top teams may be relatively accurately estimated at the beginning of the season, thus, the last rounds can be scheduled to avoid any matches between the strongest teams. Such a schedule might lead to a tanking opportunity at the end of the season with a (much) lower probability. But even if Eintracht Frankfurt had already played against RB Leipzig, it would be unfair if it turns out ex post that a team has missed qualifying for the UEFA Champions League due to winning a match months before.
Furthermore, if strong teams do not play against each other in the last rounds, some strong teams may play against an opponent with nothing to compete for anymore, while other strong teams may play against opponents still in contention for spots in UEFA club competitions or to avoid relegation. The former matches can jeopardise fairness as discussed by Csató et al., (2024) and Scelles et al., (2024). Therefore, scheduling the domestic leagues accordingly is not a good solution for the problem identified here. Consequently, it is the responsibility of UEFA to act if it is really committed to its announced principles.
6 Conclusions and managerial implications
This paper has revealed that the UEFA Champions League rules are incentive incompatible again from the 2024/25 season. In particular, the allocation rules of the vacant slots caused by the titleholders and of the two European performance spots have almost led to a game with misaligned incentives in the 2023/24 German Bundesliga. We have used a formal mathematical model to uncover the underlying causes and to provide two alternative policies: (1) Reverse the order in which the vacancies caused by the titleholders are filled and the European performance sports are allocated; or (2) Fill any vacancy caused by a titleholder within its domestic league. While the first solution substantially reduces the probability of an incentive incompatible situation, the second fully eliminates the problem if the current draw restrictions are retained. Therefore, UEFA is strongly encouraged to adopt one of our recommendations in order to safeguard the integrity of its flagship tournament.